OpenStudy (anonymous):

Construct a basis of R4 that consists of the following vectors LINEAR ALGEBRA

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \\2\\3\\4\end{matrix}\right) \left(\begin{matrix}1 \\ 4\\6\\8\end{matrix}\right)\]

OpenStudy (anonymous):

and the standard basis vectors e1,e2,e3,e4 which I'm guessing means \[\left(\begin{matrix}1 \\ 0\\0\\0\end{matrix}\right)\left(\begin{matrix}0 \\ 1\\0\\0\end{matrix}\right)\left(\begin{matrix}0 \\ 0\\1\\0\end{matrix}\right)\left(\begin{matrix}0 \\ 0\\0\\1\end{matrix}\right)\]